# BOLIMES: Boruta–LIME optiMized fEature Selection for Gene Expression Classification

### v1

@@ 4.3 Classification results
IID Dataset Class Samples Top N Classification Results Time (s)
Features Acc Prec Rec F1 Score Training Preprocess Total
1 20685 6 327 114 0,955 0,962 0,955 0,955 0,395 14,439 40,015
2 20711 5 90 10 0,833 0,841 0,833 0,830 0,328 3,800 0,328
3 21050 4 310 60 0,790 0,791 0,790 0,784 0,286 11,975 16,186
4 21122 7 158 159 0,938 0,921 0,938 0,925 0,293 4,028 44,087
5 29354 3 53 10 1,000 1,000 1,000 1,000 0,271 2,541 0,271
6 30784 3 229 50 0,978 0,979 0,978 0,976 0,127 10,126 6,344
7 31312 3 498 10 0,860 0,848 0,860 0,843 0,388 21,554 0,388
8 31552 3 111 10 1,000 1,000 1,000 1,000 0,557 4,239 0,557
9 32537 7 217 95 0,796 0,693 0,796 0,736 0,299 2,478 26,388
10 33315 10 575 425 0,870 0,842 0,870 0,855 1,341 6,495 389,225
11 36895 14 76 29 0,929 0,929 0,929 0,929 0,250 3,364 6,509
12 37364 4 94 21 1,000 1,000 1,000 1,000 0,142 4,804 1,862
13 39582 6 566 556 0,825 0,833 0,825 0,824 1,564 22,414 554,762
14 39716 3 53 44 1,000 1,000 1,000 1,000 0,101 3,074 3,746
15 44077 4 226 57 0,978 0,980 0,978 0,979 0,139 6,925 8,332
Table 6: Gradient Boosting Classification Results with Top N Features
IID Dataset Class Samples Top N Classification Results Time (s)
Features Acc Prec Rec F1 Score Training Preprocess Total
1 20685 6 327 468 0,939 0,947 0,939 0,936 17,957 12,795 17,957
2 20711 5 90 38 0,889 0,921 0,889 0,892 0,664 3,832 0,664
3 21050 4 310 65 0,807 0,802 0,807 0,804 1,844 9,923 1,844
4 21122 7 158 257 0,875 0,872 0,875 0,871 5,544 1,890 5,544
5 29354 3 53 12 1,000 1,000 1,000 1,000 0,387 0,822 0,387
6 30784 3 229 39 0,957 0,957 0,957 0,957 0,747 8,048 0,747
7 31312 3 498 169 0,880 0,853 0,880 0,854 4,961 18,680 4,961
8 31552 3 111 13 0,957 0,917 0,957 0,936 0,310 2,594 0,310
9 32537 7 217 54 0,773 0,705 0,773 0,734 2,249 0,906 2,249
10 33315 10 575 351 0,896 0,865 0,896 0,877 40,149 4,819 40,149
11 36895 14 76 11 0,857 0,857 0,857 0,857 0,826 3,260 1,645
12 37364 4 94 24 0,947 0,953 0,947 0,947 0,480 3,327 0,480
13 39582 6 566 352 0,816 0,828 0,816 0,816 24,821 20,419 24,821
14 39716 3 53 24 1,000 1,000 1,000 1,000 0,299 1,468 0,299
15 44077 4 226 18 1,000 1,000 1,000 1,000 0,711 5,468 0,711
@@
The four classification algorithms reveal distinct trade-offs between predictive performance and computational efficiency when applied to high-dimensional gene expression data. For instance, on dataset E-GEOD-20685 (ID 1), SVM attained an accuracy of 95.5% with 107 selected features and a total processing time of 112.137 seconds, while Random Forest achieved a marginally higher accuracy of 97.0% but at a prohibitive total time of 16,306.229 seconds. In contrast, XGBoost matched SVM’s accuracy (95.5%) with a considerably lower training time of 0.395 seconds and an overall time of 40.015 seconds, suggesting a more balanced performance. Gradient Boosting, on the other hand, reached an accuracy of 93.9% using 468 features, with a total time of 17.957 seconds—indicating competitive speed but slightly reduced accuracy compared to the other methods.
Similar trends are observed in other datasets; for example, in E-GEOD-29354 (ID 5), all algorithms achieved perfect accuracy, yet their feature counts and processing times varied substantially. These differences underscore the critical need for an approach that not only maintains high classification performance but also minimizes computational overhead. Although Random Forest often produces high accuracy, its excessive time cost can limit its practicality in real-world applications. Conversely, XGBoost offers a compelling balance between accuracy and efficiency, making it particularly well-suited for gene expression classification tasks. Overall, these results emphasize that algorithm selection should be guided by both predictive metrics and computational feasibility, especially in scenarios involving high-dimensional data.
4.4 Discussion
@@ 5 Conclusion and future work
In conclusion, our study introduces BOLIMES, a novel feature selection algorithm that integrates the robustness of Boruta with the interpretability of LIME to tackle the challenges inherent in gene expression classification. By systematically filtering out non-informative genes and ranking the remaining ones based on their local importance, BOLIMES effectively reduces the dimensionality of genomic data while mitigating the risk of overfitting. The iterative evaluation process further refines the feature subset by identifying the optimal number of genes that maximize predictive accuracy, thereby ensuring both high classification performance and enhanced interpretability.
Looking ahead, several promising avenues exist for further advancement of this work. Future research could explore the incorporation of additional interpretability methods alongside LIME to offer a more comprehensive assessment of feature relevance. Moreover, extending BOLIMES to accommodate multi-omics data, such as proteomic and metabolomic profiles, may broaden its applicability in complex biological analyses. Integrating deep learning-based classifiers could also be investigated to further boost performance in intricate gene expression scenarios. Finally, validating the approach on larger, more diverse datasets and real-world clinical samples will be crucial to establish its robustness and generalizability in practical settings.
Acknowledgements: This study is received support from the European Union’s Horizon research and innovation program under the MSCA-SE (Marie Skłodowska-Curie Actions Staff Exchange) grant agreement 101086252; Call: HORIZON-MSCA-2021-SE-01; Project title: STARWARS.

### v2

@@ 4.3 Classification results
ID Dataset Class Method Samples Top k Classification Results Training
Features Acc Prec Rec F1 Score Time (s)
1 20685 6 Ours 327 114 0.955 0.961 0.955 0.955 40.015
SVM 85 0.804 0.839 0.804 0.793 10.932
RF 85 0.829 0.854 0.829 0.822 11.592
DT 85 0.851 0.874 0.851 0.843 11.531
2 20711 5 Ours 90 10 0.833 0.841 0.833 0.830 0.328
SVM 43 0.744 1.000 0.744 0.744 62.459
RF 43 0.811 1.000 0.811 0.811 63.819
DT 43 0.811 1.000 0.811 0.811 63.947
3 21050 4 Ours 310 60 0.790 0.791 0.790 0.784 16.183
SVM 275 0.732 0.769 0.732 0.701 9.478
RF 275 0.748 0.782 0.748 0.724 9.977
DT 275 0.752 0.796 0.752 0.720 9.622
4 21122 7 Ours 158 159 0.938 0.921 0.938 0.924 44.087
SVM 78 0.848 1.000 0.848 0.848 162.044
RF 78 0.854 1.000 0.854 0.854 164.271
DT 78 0.867 1.000 0.867 0.867 166.522
5 29354 3 Ours 53 10 1.000 1.000 1.000 1.000 0.271
SVM 35 0.906 1.000 0.906 0.906 23.871
RF 35 0.774 1.000 0.774 0.774 23.896
DT 35 0.830 1.000 0.830 0.830 23.359
6 30784 3 Ours 229 50 0.978 0.979 0.978 0.976 6.344
SVM 42 0.921 1.000 0.921 0.921 121.816
RF 42 0.913 1.000 0.913 0.913 119.124
DT 42 0.921 1.000 0.921 0.921 123.405
7 31312 3 Ours 498 10 0.860 0.848 0.860 0.843 0.388
SVM 195 0.851 0.863 0.851 0.810 7.677
RF 195 0.833 0.828 0.833 0.797 7.432
DT 195 0.831 0.850 0.831 0.790 7.762
8 31552 3 ours 111 10 1.000 1.000 1.000 1.000 0.557
SVM 44 0.856 1.000 0.856 0.856 53.782
RF 44 0.892 1.000 0.892 0.892 52.092
DT 44 0.892 1.000 0.892 0.892 53.220
9 32537 7 ours 217 95 0.795 0.693 0.795 0.736 26.388
SVM 171 0.783 1.000 0.783 0.783 282.386
RF 171 0.802 1.000 0.802 0.802 283.478
DT 171 0.770 1.000 0.770 0.770 283.282
10 33315 10 ours 575 425 0.870 0.842 0.870 0.854 389.225
SVM 2321 0.861 0.883 0.861 0.838 94.119
RF 2321 0.859 0.879 0.859 0.839 94.236
DT 2321 0.852 0.874 0.852 0.827 96.417
11 37364 4 ours 94 21 1.000 1.000 1.000 1.000 1.862
SVM 140 0.777 1.000 0.777 0.777 59.093
RF 140 0.798 1.000 0.798 0.798 59.041
DT 140 0.745 1.000 0.745 0.745 60.416
12 39582 6 ours 566 556 0.825 0.833 0.825 0.824 554.762
SVM 441 0.797 0.810 0.797 0.793 21.655
RF 441 0.804 0.816 0.804 0.799 22.184
DT 441 0.788 0.791 0.788 0.784 22.569
13 39716 3 ours 53 44 1.000 1.000 1.000 1.000 3.746
SVM 118 0.897 1.000 0.887 0.887 23.631
RF 118 0.925 1.000 0.925 0.925 22.742
DT 118 0.943 1.000 0.943 0.943 24.227
14 44077 4 ours 226 57 0.978 0.980 0.978 0.978 8.332
SVM 23 0.978 1.000 0.978 0.978 132.205
RF 23 0.987 1.000 0.987 0.987 133.960
DT 23 0.991 1.000 0.991 0.991 133.546
Table 6: Gradient Boosting Classification Results with Top $k$ Features
ID Dataset Class Method Samples Top k Classification Results Training
Features Acc Prec Rec F1 Score Time (s)
1 20685 6 ours 327 468 0.939 0.947 0.939 0.936 17.957
SVM 85 0.795 0.820 0.795 0.788 53.840
RF 85 0.856 0.876 0.856 0.855 53.966
DT 85 0.844 0.865 0.844 0.834 53.381
2 20711 5 ours 90 38 0.889 0.921 0.889 0.892 0.664
SVM 43 0.733 1.000 0.733 0.733 91.394
RF 43 0.711 1.000 0.711 0.711 90.162
DT 43 0.711 1.000 0.711 0.711 89.505
3 21050 4 ours 310 65 0.806 0.802 0.806 0.804 1.844
SVM 275 0.703 0.733 0.703 0.685 72.745
RF 275 0.735 0.745 0.735 0.720 78.348
DT 275 0.758 0.792 0.758 0.742 76.235
4 21122 7 ours 158 257 0.875 0.872 0.875 0.871 5.544
SVM 78 0.804 1.000 0.804 0.804 511.805
RF 78 0.835 1.000 0.835 0.835 507.856
DT 78 0.861 1.000 0.861 0.861 513.399
5 29354 3 ours 53 12 1.000 1.000 1.000 1.000 0.387
SVM 35 0.849 1.000 0.849 0.849 22.807
RF 35 0.717 1.000 0.717 0.717 20.737
DT 35 0.811 1.000 0.811 0.811 20.797
6 30784 3 ours 229 39 0.957 0.957 0.957 0.957 0.747
SVM 42 0.913 1.000 0.913 0.913 272.929
RF 42 0.900 1.000 0.900 0.900 269.238
DT 42 0.943 1.000 0.943 0.943 273.324
7 31312 3 ours 498 169 0.880 0.853 0.880 0.854 4.961
SVM 195 0.853 0.840 0.853 0.830 89.338
RF 195 0.853 0.828 0.853 0.824 90.304
DT 195 0.825 0.805 0.825 0.795 90.659
8 31552 3 ours 111 13 0.957 0.917 0.957 0.936 0.310
SVM 44 0.802 1.000 0.802 0.802 77.917
RF 44 0.883 1.000 0.883 0.883 74.657
DT 44 0.892 1.000 0.892 0.892 70.156
9 32537 7 ours 217 54 0.773 0.705 0.773 0.734 2.249
SVM 171 0.747 1.000 0.747 0.747 1942.324
RF 171 0.742 1.000 0.742 0.742 1928.460
DT 171 0.747 1.000 0.747 0.747 1943.035
10 33315 10 ours 575 351 0.896 0.865 0.896 0.877 40.149
SVM 2321 0.833 0.847 0.833 0.822 4118.864
RF 2321 0.856 0.873 0.856 0.839 4055.604
DT 2321 0.840 0.852 0.840 0.824 3923.718
11 37364 4 ours 94 24 0.947 0.953 0.947 0.947 0.480
SVM 140 0.787 1.000 0.787 0.787 173.044
RF 140 0.745 1.000 0.745 0.745 172.476
DT 140 0.745 1.000 0.745 0.745 172.407
12 39582 6 ours 566 352 0.816 0.828 0.816 0.816 24.821
SVM 441 0.802 0.819 0.802 0.799 465.967
RF 441 0.813 0.819 0.813 0.806 465.414
DT 441 0.776 0.794 0.776 0.768 470.792
13 39716 3 ours 53 24 1.000 1.000 1.000 1.000 0.299
SVM 118 0.906 1.000 0.906 0.906 38.331
RF 118 0.868 1.000 0.868 0.868 34.458
DT 118 0.811 1.000 0.811 0.811 38.443
14 44077 4 ours 226 18 1.000 1.000 1.000 1.000 0.711
SVM 23 0.973 1.000 0.973 0.973 220.643
RF 23 0.982 1.000 0.982 0.982 210.123
DT 23 0.987 1.000 0.987 0.987 224.643
@@
The four classification algorithms reveal distinct trade-offs between predictive performance and computational efficiency when applied to high-dimensional gene expression data. For instance, on dataset E-GEOD-20685 (ID 1), SVM attained an accuracy of 95.5% with 107 selected features and a total processing time of 112.137 seconds, while Random Forest achieved a marginally higher accuracy of 97.0% but at a prohibitive total time of 16,306.229 seconds. In contrast, XGBoost matched SVM’s accuracy (95.5%) with a considerably lower training time of 0.395 seconds and an overall time of 40.015 seconds, suggesting a more balanced performance. Gradient Boosting, on the other hand, reached an accuracy of 93.9% using 468 features, with a total time of 17.957 seconds—indicating competitive speed but slightly reduced accuracy compared to the other methods. Similar trends are observed in other datasets; for example, in E-GEOD-29354 (ID 5), all algorithms achieved perfect accuracy, yet their feature counts and processing times varied substantially. These differences underscore the critical need for an approach that not only maintains high classification performance but also minimizes computational overhead. Although Random Forest often produces high accuracy, its excessive time cost can limit its practicality in real-world applications. Conversely, XGBoost offers a compelling balance between accuracy and efficiency, making it particularly well-suited for gene expression classification tasks.
In general, the results of our proposed method outperformed feature selection techniques in terms of accuracy. Moreover, the computational time for the training process was significantly reduced, as our approach identified fewer but more crucial and promising features compared to previous studies. This reduction in the number of features also led to a faster training time. In some cases, although more features were selected, the accuracy achieved was considerably higher than that of other solutions
@@ 5 Conclusion and future work
This study presents BOLIMES, a novel feature selection algorithm that combines the robustness of Boruta with the interpretability of LIME to address the challenges in gene expression classification. The proposed method has demonstrated superior performance, achieving higher accuracy compared to alternative solutions, while significantly reducing training time. Furthermore, we have identified the optimal number of features necessary to achieve the best possible accuracy. Looking ahead, future work could focus on integrating additional interpretability methods with LIME for a more comprehensive feature relevance assessment. Extending BOLIMES to multi-omics data, including proteomics and metabolomics, could enhance its applicability. Investigating deep learning-based classifiers and validating the approach on larger, diverse datasets and clinical samples would be essential for improving performance and ensuring robustness in real-world applications.
Acknowledgements: Thanh MA and Thanh-Nghi DO are received support from the European Union’s Horizon research and innovation program under the MSCA-SE (Marie Skłodowska-Curie Actions Staff Exchange) grant agreement 101086252; Call: HORIZON-MSCA-2021-SE-01; Project title: STARWARS.